Two angles are complementary. The sum of the measure of the first angle and one fourth the second angle is 75.75 degrees. Find the measures of the angles.
*what is the measure of the smaller angle?
*what is the measure of the other angle?

Answer by nerdybill(6962)

Two angles are complementary. The sum of the measure of the first angle and one fourth the second angle is 75.75 degrees. Find the measures of the angles.
Let x = first angle
then
90-x = complementary of first angle
.
From: "The sum of the measure of the first angle and one fourth the second angle is 75.75 degrees."
x + (1/4)(90-x) = 75.75
x + 22.5-.25x = 75.75
.75x + 22.5 = 75.75
.75x = 53.25
x = 53.25/.75
x = 71 degrees
.
Complementary angle:
90-x = 90-71 = 19 degrees
.
*what is the measure of the smaller angle? 19 degrees
*what is the measure of the other angle? 71 degrees